# Umberto Collodel Central Bank of Malta

### v1

@@ 4.4 Simple Benchmark Comparison
Our LLM-based framework captures heterogeneous market reactions to ECB press conferences with substantial accuracy. Correlations between the simulated disagreement and actual post-announcement OIS volatility reach roughly 0.5 for medium- and long-term tenors, with performance further enhanced under the LLM-as-a-Judge framework at early iterations. These results remain stable across training and test datasets, indicating that the model captures fundamental communication-market relationships rather than memorizing patterns.
To contextualize these findings, we next compare the LLM-based correlations with those obtained using a simple textual benchmark, highlighting the added value of modeling heterogeneous market interpretations.
@@
Specifically, we examine correlations between post-ECB press conference OIS volatility and the Flesch–Kincaid readability score of the statements (Figure 10). [^28] The Flesch–Kincaid readability score is a standard measure of text readability that estimates the U.S. school grade level required to understand a given text. Higher scores indicate more complex language, while lower scores indicate easier-to-read text. In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns than simpler ones.
@@
Our own correlations are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively. Taken together, these findings suggest that readability and complexity are indeed part of the story: more complex communication can amplify market uncertainty. Nevertheless, this dimension is only one facet of the broader interpretive process. By construction, readability indices cannot capture policy content, such as the magnitude of the first-moment surprise or communication shifts like the abandonment of forward guidance, so their explanatory power remains limited. [^29] For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock. Yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility. In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
@@ 5 Conclusions
Our key findings advance both theoretical understanding and practical application in monetary economics. First, even a naive zero-shot approach achieves substantial correlations (~0.5) with actual market disagreement, particularly for medium and long-term maturities. This suggests LLMs possess intrinsic capabilities to decode central bank communication in economically meaningful ways. Second, contrary to other studies employing LLMs for macroeconomic forecasting (i.e. Hansen et al., (2024)), few-shot prompting with historical context does not improve performance. This highlights that shared examples can induce excessive anchoring when employed in volatily simulation rather than pure point estimates. Third, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization. The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
These results have significant practical implications. For central banks, our framework offers a powerful ex-ante tool to evaluate communication strategies and identify potentially market-disrupting language before release. For market participants, it provides insights into the drivers of post-announcement volatility. For researchers, it opens the ”black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies. The methodology is readily adaptable to other central banks and policy contexts.
Despite these contributions, a few limitations remain. The replicability of our results is partly constrained by the versioning and closed-source nature of contemporary LLMs, meaning that future iterations of the same model may yield slightly different outputs. A natural way forward would be to employ open-source models, which would allow for greater control, transparency, and long-term reproducibility. [^30] See Rogers and Luccioni, (2023) for a thorough discussion. Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication—or analyses thereof—were included in the model’s pretraining data cannot be completely dismissed. In our setup, we explicitly instruct the model to use only information available up to the date of each press conference; however, this constraint affects its responses, not its underlying parameters. Nevertheless, we consider the risk of substantial contamination to be limited, as discussions of post-ECB press conference volatility are relatively rare in the public domain, with only a small number of academic papers addressing the topic. Future research should focus on tackling these issues. Moreover, we consider these results a lower bound: testing with domain-specific LLMs or fine-tuned models could shed light on the trade-off between interpretive diversity and specialized expertise, potentially improving both the accuracy and the behavioral realism of simulated market reactions.
@@
Appendix

### v2

@@ 5.1 Simple Benchmark Comparison
A critical question to justify our approach is whether sophisticated LLM-based agent modeling offers genuine advantages over simple textual measures, or whether basic complexity metrics can achieve similar predictive power at much lower computational cost. To address this fundamental methodological question, we compare the LLM-based correlations with a simple textual benchmark: the Flesch-Kincaid readability score. The Flesch-Kincaid score is a standard measure of text readability that estimates the U.S. school grade level required to understand a given text, with higher scores indicating more complex language.
In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns compared to simpler ones. Our analysis tests whether this relationship holds for ECB communications and how it compares to our LLM-based approach.
@@
The correlations between Flesch-Kincaid scores and post-conference OIS volatility are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively (Figure 10). These results suggest that readability and complexity might be part of the story: more complex communication can amplify market uncertainty. Nevertheless, this dimension captures only one small facet of the broader interpretive process. By construction, readability indices cannot capture policy content, such as the magnitude of the monetary policy surprise, or strategy shifts, like the abandonment of forward guidance, so their explanatory power remains limited. For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock, yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility. In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
5.2 Model Architecture Robustness
To assess consistency across different LLM architectures, we use the Intraclass Correlation Coefficient (ICC), a reliability metric commonly used to evaluate agreement between multiple measurement methods applied to the same subjects. In this case, ICC measures the proportion of total variance attributable to true differences between conferences relative to total observed variance, including measurement error.
Formally,
$\text{ICC}_{t}=\frac{\sigma^{2}_{\text{conferences}}}{\sigma^{2}_{\text{conferences}}+\sigma^{2}_{\text{methods}}}$ (1)
where $\sigma^{2}_{\text{conferences}}$ is the variance of disagreement across ECB conferences (signal), and $\sigma^{2}_{\text{methods}}$ is the variance across different models or prompt specifications (noise). The subscript $t$ corresponds to the OIS tenor. ICC values range from 0 to 1, with higher values indicating that observed variance reflects genuine conference differences rather than methodological noise. Following established conventions, ICC values below 0.40 are considered poor, 0.40-0.59 fair, 0.60-0.74 good, and 0.75-1.00 excellent (Cicchetti,, 1994).
We assess consistency across three distinct LLM architectures: Gemini, ChatGPT, and Claude. [^20] Specifically, we use Gemini 2.5-Flash, ChatGPT-5-mini, and Claude-4-Sonnet. These models were selected to capture diversity in training data, design philosophy, and reasoning capabilities. For each ECB conference, we generate synthetic disagreement measures, compute correlations with market-based values, and evaluate cross-model agreement using ICCs.
Figure 11: Average ICC across Model Variation by Tenor and Ensemble Model Correlation with Market-based disagreement
Note: Blue bars show the intraclass correlation coefficient for each tenor calculated using three distinct LLM architectures (Gemini 2.5 Flash, ChatGPT 5-mini, Claude 4 Sonnet). Higher ICC values indicate greater measurement stability. Red bars show the Spearman correlation between the ensemble of LLM outputs, a simple average, and market-based disagreement.
The results show fair agreement among models: 3-month tenor (ICC = 0.499), 2-year tenor (ICC = 0.466), and 10-year tenor (ICC = 0.408). While these ICC values are acceptable according to conventional thresholds, they fall in the lower-moderate range and warrant some caution in interpretation. Nevertheless, we investigate the matter further and find that each individual model achieves correlations of approximately 0.5 across all tenors with market-based uncertainty measures (Figure 16). Therefore, we hypothesize that different models may capture complementary aspects of ECB communication, rather than indicating noisy measurements. To verify this hypothesis, we compute an ensemble measure as the simple average of all three models’ output. The ensemble results strongly support our interpretation: the combined measure achieves even higher correlations than any individual model (Figure 11).
These findings support two key conclusions: while individual LLMs are not perfectly interchangeable, their systematic differences capture distinct but relevant aspects of communication interpretation; ensemble methods can harness this diversity to provide more complete measures of market-relevant uncertainty than any single model alone.
5.3 Prompt Specification Robustness
LLMs are particularly sensitive to prompt variations, with “performance differences of up to 76 percentage points for subtle formatting changes” (Sclar et al.,, 2024). To assess this vulnerability in our context, we generate 15 prompt perturbations of our baseline specification, including 10 minor and 5 medium variations. Minor variations substitute semantically equivalent alternatives and reorder prompt components, while medium variations employ substantively different language to describe trader characteristics and market conditions while maintaining the overall framework. Table 1 shows a representative set of examples. [^21] Full list in Table 2 in Appendix A. We test these variations on 30 randomly selected conferences, computing ICCs to measure reliability across prompt specifications. [^22] Given the large computational resources required to run all perturbations on all conferences (283 conferences × 15 perturbations = 4,245 API calls), we select a random subsample of 30 conferences to balance statistical power with computational feasibility.
Table 1: Examples of Prompt Perturbations
Perturbation Type Intensity Example Modification Methodological Purpose Agent Characteristics Minor Changed “Risk Aversion: High/Medium/Low” to “Risk Tolerance: Conservative/Moderate/Aggressive” Test sensitivity to terminology while preserving semantic meaning Medium Replaced behavioral biases list: “Confirmation Bias, Overconfidence, Anchoring” $\rightarrow$ “Home Bias, Momentum Trading, Disposition Effect” Evaluate robustness to different but equivalent behavioral frameworks Task Instructions Minor “Provide an expected rate direction” $\rightarrow$ “Indicate your directional forecast” Assess impact of instruction phrasing variations Medium Added explicit confidence calibration: “Confidence (0-100%) reflecting forecast certainty” $\rightarrow$ “Confidence (0-100%) based on information clarity and personal conviction” Test robustness to enhanced guidance specificity Output Format Minor Changed table headers: “New Expected Rate (%)” $\rightarrow$ “Predicted Rate (%)” Verify stability across formatting variations Medium Modified output structure: Single table $\rightarrow$ Separate tables per tenor with additional reasoning column Examine sensitivity to output complexity changes Agent Population Minor “30 individual traders” $\rightarrow$ “30 market participants” Assess sensitivity to agent terminology Medium Changed agent metaphor: “traders” $\rightarrow$ “institutional investors with varying mandates” Test robustness to different agent conceptualizations
Reliability follows a clear maturity hierarchy: the 3-month tenor exhibits the highest overall reliability (average ICC = 0.655), followed by the 2-year (ICC = 0.595) and 10-year (ICC = 0.548) tenors (Figure 12).
Figure 12: Average ICC across Prompt Variations by Tenor
Note: Average ICC values across all prompt variations (10 minor and 5 medium) for each tenor. Higher ICC values indicate greater measurement stability. Variations are based on prompt 2 and obtained by calling Gemini 2.5-Flash. Dashed horizontal lines indicate 0.4 (acceptable ICC), 0.6 (good ICC), and 0.75 (excellent ICC) thresholds based on Cicchetti, (1994).
The 10-year tenor exhibits greater prompt sensitivity, with medium variations substantially outperforming minor ones (0.648 vs 0.448) (Figure 13). This pattern partly reflects a data limitation: unlike shorter maturities where market data extend back to 2005, daily 10-year OIS data are available only from 2010, reducing both sample size and coverage of high-uncertainty periods. Nevertheless, all tenors maintain ICC values above conventional reliability thresholds.
Figure 13: Minor vs Medium Variations ICC by Tenor
Note: Comparison of ICC between minor and medium prompt variations for each tenor. Higher ICC values indicate greater measurement stability. Variations are based on prompt 2 and obtained by calling Gemini 2.5-Flash. Dashed horizontal lines indicate 0.4 (acceptable ICC), 0.6 (good ICC), and 0.75 (excellent ICC) thresholds based on Cicchetti, (1994).
The perturbation analysis demonstrates substantial robustness across all tenor-specification combinations. ICC values consistently exceed acceptable reliability thresholds (0.4), frequently surpass good reliability benchmarks (0.6), and occasionally approach excellent standards (0.75). These results indicate that conference-specific variation systematically dominates prompt-specific noise.
@@ 6 Conclusions
Our key findings advance both theoretical understanding and practical application in monetary economics. First, even a naive zero-shot approach achieves substantial correlations (roughly 0.5) with actual market disagreement, particularly for medium and long-term maturities. This suggests LLMs can effectively decode central bank communication in economically meaningful ways. Second, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization. The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
These results have significant practical implications. For central banks, our framework offers a powerful ex-ante tool to evaluate communication strategies and identify potentially market-disrupting language before release. For market participants, it provides insights into the drivers of post-announcement volatility. For researchers, it opens the “black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies and rule-based models. The methodology is flexible and readily adaptable to other central banks and policy contexts.
Despite these contributions, some limitations remain. The closed-source nature of leading LLMs constrains replicability, as future model versions may yield slightly different outputs. While we mitigate this problem through multi-model robustness testing, future research could explore open-source alternatives for greater transparency and reproducibility. Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication, or analyses thereof, were included in the model’s pretraining data cannot be completely dismissed. We explicitly instruct the model to use only information available up to each conference date; however, this constraint affects responses, not underlying parameters. Nevertheless, we consider substantial contamination risk limited given that discussions of post-ECB press conference volatility are rare in the public domain. Finally, domain-specific or fine-tuned models could potentially enhance performance by balancing interpretive diversity with specialized financial expertise, though this remains to be tested.
@@
Appendix A: Additional Sensitivity Checks And Robustness Details
6.1 Different Correlation Measures
To ensure the robustness of our main findings, we conduct a supplementary analysis verifying that the observed relationship between synthetic and market-based disagreement is not dependent on the choice of the correlation metric. Our baseline analysis employs the Spearman rank correlation coefficient ($\rho$), which is well-suited for this context as it captures monotonic relationships and is robust to outliers and non-linearities.
For a more comprehensive validation, we also consider two additional measures: the Pearson linear correlation coefficient ($r$) and Kendall’s rank correlation coefficient ($\tau$):
  • Pearson’s $r$ quantifies the strength of a strictly linear relationship between two variables. Although more sensitive to outliers, it allows us to check whether the relationship could plausibly be explained within a linear framework.
  • Kendall’s $\tau$ provides another non-parametric, rank-based alternative that evaluates concordance between two rankings. It is often preferred in smaller samples or in the presence of tied ranks, offering an additional robustness check alongside Spearman’s $\rho$.
To assess the precision of these estimates, we compute 95% confidence intervals via a non-parametric bootstrap with 5,000 replications. Figure 14 presents the results.
Figure 14: Comparison of Correlation Measures Between Synthetic and Market-Based Disagreement.
Note: 95% confidence intervals computed via a non-parametric bootstrap with 5,000 replications.
Across all three metrics, we find a statistically and economically significant positive association between our LLM-generated disagreement measure and market volatility. The 2-year tenor exhibits the strongest association in the rank-based measures (Spearman’s $\rho=0.53$; Kendall’s $\tau=0.37$), highlighting the simulation’s effectiveness at capturing uncertainty around the medium-term monetary policy path.
Importantly, the bootstrapped 95% confidence intervals for all coefficients across all tenors comfortably exclude zero, underscoring statistical significance. While the magnitudes vary across correlation metrics, consistent with their differing assumptions, the overall conclusion remains unchanged. The alignment of results across linear and rank-based approaches demonstrates that the link we identify is a robust and stable feature of the data.
6.2 Sensitivity to Market Volatility Window
To ascertain whether our baseline findings are contingent upon the specific 3-day window used to measure market volatility, we conduct a robustness check using alternative 1-day and 2-day post-conference windows. The results shown in Figure 15 confirm that the positive and statistically significant relationship between our synthetic disagreement measure and market volatility is robust across all tested specifications.
Figure 15: Correlation between Synthetic Disagreement and Market-Based Calculated Over Different Windows
Note: Market- based definitions are equal to the min-max range in the day of the press conference, the average of the day of the conference and the following day, and within a three days period (baseline). 95% confidence intervals computed via a non-parametric bootstrap with 5,000 replications.
The analysis reveals two key patterns. First, for the 2-year tenor, the Spearman correlation is both high and remarkably stable (in the ballpark of 0.5). This stability reinforces the conclusion that the simulation is particularly adept at capturing uncertainty over the medium-term policy path, which appears to be the dominant factor in the days immediately following the communication event. Second, for the 10-year tenor, the correlation is strongest in the 1-day window (0.49) before declining slightly in the 2- and 3-day windows (0.44). This suggests our LLM-based framework is exceptionally effective at capturing the immediate, pure market interpretation of long-term policy signals. The subsequent modest decline in correlation may reflect the re-emergence of other long-term macroeconomic factors that influence the 10-year rate beyond the initial communication shock. Nevertheless, confidence intervals are largely overlapping. Consistent with our main findings, the 3-month tenor exhibits a statistically significant but weaker correlation across all windows.
6.3 Model perturbation
Across all three major language models tested, ChatGPT, Claude, and Gemini, the 2-year tenor synthetic value consistently achieves the highest correlations with actual market volatility, with Claude delivering the strongest performance at 0.586, followed by Gemini at 0.526 and ChatGPT at 0.440. This pattern reinforces our earlier findings that medium-term maturities, which primarily capture monetary policy expectations shaped by ECB forward guidance, are most amenable to LLM-based interpretation. The 10-year tenor shows more variable performance across models, with Claude again leading (0.563) but with a notably wider spread between the best and worst performing models (0.563 vs 0.287 for ChatGPT), suggesting that longer-term yield dynamics present greater challenges for consistent LLM interpretation. At the short end, the 3-month tenor exhibits the most uniform performance across models, with correlations ranging from 0.389 to 0.467, reflecting the inherent constraints of modeling near-term rates during periods of policy predictability. Importantly, Claude’s superior performance across all tenors may reflect architectural differences in processing complex financial text, though the substantial correlations achieved by all three models (consistently above 0.28) demonstrate that the core phenomenon, LLMs’ ability to capture interpretive heterogeneity in response to central bank communication, is not model-specific but rather represents a fundamental capability of this class of models. [^23] Initially, we had scraped some conferences included in the monetary policy section of the ECB website that were not actually about monetary policy (due to the ECB’s classification error). When running all the LLMs on those conferences, Claude was the only one that did not create the usual table, instead responding that those texts were not about monetary policy decisions. This demonstrates Claude’s superior content validation capabilities and suggests higher reliability in distinguishing relevant monetary policy communication from unrelated content.
Figure 16: Individual Model Correlations with Market-Based Disagreement by Tenor
Note: Spearman correlation coefficients between synthetic disagreement measures generated by individual LLM models (Gemini 2.5-Flash, ChatGPT 5-mini, and Claude Sonnet 4) and market-based post-conference OIS volatility across three tenors (3-month, 2-year, and 10-year). Each model generates synthetic disagreement by simulating 30 heterogeneous traders interpreting ECB press conference transcripts. Market-based disagreement is measured as the average min-max range of OIS rates in the 3 days following ECB press conferences. Sample covers 283 ECB Governing Council meetings from June 1998 to April 2025.
6.4 Prompt perturbation
Table 2: List of Prompt Variations
Variation Type Key Changes Modified Section
1 Minor Individual traders → distinct market participants Context
2 Minor Trading decision → investment choice Context
3 Minor Risk aversion order: High/Medium/Low → Low/Medium/High Trader Characteristics
4 Minor Interpretation of conference → analysis of meeting Context
5 Minor Markdown table → structured table format Guidelines
6 Minor Confidence score → certainty level Task
7 Minor Monetary policy decisions → interest rate policies Context
8 Minor Maximize profit → optimize returns Context
9 Minor Reordered behavioral biases (Anchoring first) Trader Characteristics
10 Minor Press conference → policy meeting throughout Multiple sections
11 Medium Formal academic language; market agents Context
12 Medium Behavioral lenses; psychological profiles emphasis Context
13 Medium Risk tolerance: Conservative/Moderate/Aggressive Trader Characteristics & Context
14 Medium Policy signals; forward guidance; processing approach Context & Characteristics
15 Medium Position-taking behavior; different output headers Task & Output
Appendix B: LLM-as-Judge Implementation Details and Results